# Sufficient and Necessary Barrier-like Conditions for Safety and Reach-avoid Verification of Stochastic Discrete-time Systems

### v2

@@ Abstract
In this paper, we examine necessary and sufficient barrier-like conditions for infinite-horizon safety verification and reach-avoid verification of stochastic discrete-time systems, derived through a relaxation of Bellman equations. Unlike previous methods focused on barrier-like conditions that primarily address sufficiency, our work rigorously integrates both necessity and sufficiency for properties pertaining to infinite time. Safety verification aims to certify the satisfaction of the safety property, which stipulates that the probability of the system, starting from a specified initial state, remaining within a safe set always is greater than or equal to a specified lower bound. A necessary and sufficient barrier-like condition is formulated for safety verification. In contrast, reach-avoid verification extends beyond safety to include reachability, seeking to certify the satisfaction of the reach-avoid property. It requires that the probability of the system, starting from a specified initial state, reaching a target set eventually while remaining within a safe set until the first hit of the target, is greater than or equal to a specified lower bound. Two necessary and sufficient barrier-like conditions are formulated under certain assumptions.
@@ I Introduction
Temporal verification is crucial in modern systems analysis, particularly in complex systems where temporal behavior is of paramount importance [17]. It involves rigorously examining a system’s adherence to temporal properties, including safety and reach-avoid guarantees, to ensure desired outcomes and avoid undesirable events. Formal methods like model checking [5] and theorem proving [13] are indispensable tools in this process, allowing for precise and comprehensive analysis of temporal specifications.
Over the past two decades, barrier certificates have emerged as a powerful tool for safety and reach-avoid verification of dynamical systems. They provide a Lyapunov-like assurance regarding system behavior, with the mere existence of a barrier certificate being sufficient to establish the satisfiability of safety and reach-avoid specifications, as demonstrated in [17]. This approach simplifies the verification process and offers a formal mathematical framework for ensuring safety and correctness of a system without the need to explicitly evolve it over time. Especially, with the advancement in polynomial optimization, particularly sum-of-squares polynomial optimization, the problem of finding barrier certificates can be addressed through convex optimization, especially when the system of interest is polynomial, further motivating the development of barrier certificate-based methods. Barrier certificates were initially proposed for deterministic systems as a popular formal approach to safety verification in [14]. Subsequent efforts have focused on adapting and enhancing barrier functions, as well as broadening their applications [9, 2, 8, 3]. However, many real-world applications are susceptible to stochastic disturbances and are thus modeled as stochastic systems. In the continuous-time stochastic setting, safety verification over the infinite time horizon via barrier certificates was introduced alongside its deterministic counterpart in [15]. Based on Ville’s Inequality [23] and a stopped process, [15] developed a non-negative barrier function and established a sufficient condition for safety verification over the infinite time horizon, i.e., to certify upper bounds of probabilities of eventually entering an unsafe region from specific initial states, while ensuring that the system remains within the interior of a state-constrained set until the first encounter with the unsafe set. Subsequently, drawing inspiration from [10], [19] formulated a sufficient barrier-like condition for upper-bounding the probability of entering an unsafe region from certain initial states within finite time frames, while maintaining the system’s presence within the interior of a state-constrained set until the initial contact with the unsafe set. The systems in [19] involve both continuous-time and discrete-time ones. Especially, when the state-constrained set is a robust invariant (i.e., systems initialized within it remain within it under all disturbances) and the unsafe set is a subset of the invariant set, sufficient barrier-like conditions for safety verification of stochastic discrete-time systems were studied in [4, 31]. Another commonly studied safety property is related to set invariance. It is to justify lower bounds of safety probabilities over either the infinite time horizon (i.e., the system stays within a specified safe set for all time) or finite time horizons (i.e., the system stays within a given safe set during a specified finite time period) [1]. In other words, it is justifying upper bounds of exit probabilities that the system exits a specified safe set either eventually or within a bounded time horizon. Correspondingly, sufficient barrier-like conditions were developed for such safety verification over finite time horizons (e.g., [21, 11, 7, 6]) and the infinite time horizon (e.g., [29]). Afterwards, control barrier functions for synthesizing controllers to guarantee safety were explored in [20, 24]. Though finite-time verification is adequate for systems with bounded operational horizons, this work prioritizes infinite-time safety guarantees—essential for scenarios requiring robustness against indefinite or unpredictable mission lifetimes. It is worth noting here that the proposed method can also be applied to the safety verification scenario over the infinite time horizon in [15]. As to the reach-avoid verification, more recently, a new sufficient barrier-like condition was proposed in [27] for reach-avoid analysis of stochastic discrete-time dynamical systems over the infinite time horizon and later, extended to stochastic continuous-time dynamical systems [28]. This condition was constructed by relaxing a set of equations whose solution characterizes the exact reach-avoid probability of eventually entering a desired target set from an initial state while adhering to safety constraints. Another barrier-like function, termed reach-avoid supermartingales, was proposed to guarantee reach-avoid specifications as well as facilitate controllers synthesis for stochastic discrete-time systems in [32, 33]. This framework assumes the system evolves within a compact robust invariant set, with both the unsafe set and target set confined as subsets of this invariant domain. This strong assumption constrains the applicability of the framework, as many systems do not possess compact robust invariant sets, as discussed in [29]. These barrier-like conditions aim to lower bound reach-avoid probabilities in [27, 32, 33, 28]. On the other hand, converse theorems for barrier certificates, which focus on the existence of barrier certificates, have significantly contributed to elucidating how safety and reach-avoid criteria can be represented by barrier certificates. Consequently, these concepts have garnered growing interest since the inception of barrier certificates, and have been investigated in [16, 17, 25, 18, 12]. Nonetheless, there exists a paucity of research exploring the existence of barrier certificates for stochastic dynamical systems. This work aims to address this void.
@@
By relaxing Bellman equations, this paper derives necessary and sufficient barrier-like conditions for verifying safety and reach-avoid properties in stochastic discrete-time systems over infinite-time horizons. The safety verification process involves assessing whether the safety probability that a system, starting from an initial state, will stay within a safe set for all time is greater than or equal to a specified threshold. By relaxing a Bellman equation, one of whose solutions characterizes the exact safety probability, we construct a necessary and sufficient barrier-like condition for safety verification. On the other hand, the reach-avoid verification concerns verifying whether the reach-avoid probability that the system, starting from an initial state, will enter a target set eventually while avoiding unsafe sets before hitting the target, is greater than or equal to a specified threshold. We consider two cases for the reach-avoid verification. In the first case, we assume that the system will either enter the target set or leave the safe set in finite time almost surely. Under this context, by relaxing a Bellman equation, which possesses a unique bounded solution that characterizes the exact reach-avoid probability, we construct a necessary and sufficient barrier-like condition for the reach-avoid verification. In the second case, we assume that the specified threshold is strictly smaller than the exact reach-avoid probability. Under this context, by relaxing a Bellman equation featuring a unique bounded solution that provides a lower bound of the exact reach-avoid probability, we construct a necessary and sufficient barrier-like condition for the reach-avoid verification.
@@
This paper is structured as follows: Section II formulates the stochastic discrete-time systems of interest and the safety and reach-avoid verification problems. Section III presents necessary and sufficient barrier-like conditions for safety verification and their derivation. Afterward, Section IV presents necessary and sufficient barrier-like conditions for reach-avoid verification and their derivation. Finally, we conclude this paper in Section V.
@@ II Preliminaries
We start the exposition by a formal introduction of stochastic discrete-time systems and safety/reach-avoid verification problems of interest. Before posing the problem studied, let me introduce some basic notions used throughout this paper: $\mathbb{R}$ denotes the set of real values; $\mathbb{N}$ denotes the set of nonnegative integers; $\mathbb{N}_{\leq k}$ is the set of non-negative integers being less than or equal to $k$; $\mathbb{N}_{\geq k}$ is the set of non-negative integers being larger than or equal to $k$; for sets $\Delta_{1}$ and $\Delta_{2}$, $\Delta_{1}\setminus\Delta_{2}$ denotes the difference of sets $\Delta_{1}$ and $\Delta_{2}$, which is the set of all elements in $\Delta_{1}$ that are not in $\Delta_{2}$; $1_{A}(\bm{x})$ denotes the indicator function in the set $A$, where, if $\bm{x}\in A$, then $11_{A}(\bm{x})=$ 1 and if $\bm{x}\notin A$, $01_{A}(\bm{x})=$ 0.
@@ II-A Problem Statement
A disturbance signal $\pi$ is a sample path of the stochastic process $\{\bm{\theta}(i):\Theta\rightarrow\Theta,i\in\mathbb{N}\}$, which is defined on the canonical sample space $\Theta^{\infty}$, endowed with its product topology $\mathcal{B}(\Theta^{\infty})$, with the probability measure $\mathbb{P}_{\pi}:=\mathbb{P}_{\bm{\theta}}^{\infty}$. The expectation associated with the probability measure $\mathbb{P}_{\pi}$ is denoted by $\mathbb{E}_{\pi}[\cdot]$.
@@
The safety and reach-avoid verification for the system (1) over the infinite time horizon are defined below.
@@
Given a safe set $\mathcal{X}\subseteq\mathbb{R}^{n}$, an initial state $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$, a target set $\mathcal{X}_{r}\subseteq\mathcal{X}$, and a lower bound $\epsilon_{1}\in[0,1]$, the safety verification aims to certify that the safety probability $\mathbb{P}_{\pi}(S_{\bm{x}_{0}})$, which denotes the probability that the system (1), starting from the initial state $\bm{x}_{0}$, will stay within the safe set $\mathcal{X}$ for all time, is greater than or equal to $\epsilon_{1}$, i.e.,
$\mathbb{P}_{\pi}(S_{\bm{x}_{0}})\geq\epsilon_{1}.$
@@
In the sequel, we will formulate necessary and sufficient barrier-like conditions for certifying $\epsilon_{1}\leq\mathbb{P}_{\pi}(S_{\bm{x}_{0}})$. It is worth noting here that the proposed method can also be used to construct necessary and sufficient conditions for the safety verification scenario in [15], which involves certifying upper bounds of the probability that the system eventually enters unsafe sets from an initial state while adhering to state-constrained sets. Please refer to Remark 2 in Subsection IV-A. In contrast, under certain assumptions, we will formulate necessary and sufficient barrier-like conditions for certifying $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})\geq\epsilon_{2}$.
@@ III Safety Verification
This section will introduce necessary and sufficient barrier-like conditions for certifying lower bounds in safety verification and will detail their construction process. The construction involves constructing and relaxing a Bellman equation, one of whose solutions characterizes the exact safety probability $\mathbb{P}_{\pi}(S_{\bm{x}})$ for $\bm{x}\in\mathbb{R}^{n}$. The Bellman equation is derived from a value function.
Let’s start with the value function $V(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$, which is able to characterize the exact safety probability $\mathbb{P}_{\pi}(S_{\bm{x}})$ for $\bm{x}\in\mathbb{R}^{n}$,
@@
Clearly, $\mathbb{E}_{\pi}[1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(0))]=1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{x})$. In addition, since
$\begin{split}\mathbb{E}_{\pi}[\prod_{j=0}^{i-1}1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(i))]=\mathbb{P}_{\pi}(\wedge_{j=1}^{i-1}[\bm{\phi}_{\pi}^{\bm{x}}(j)\in\mathcal{X}]\wedge[\bm{\phi}_{\pi}^{\bm{x}}(i)\in\mathbb{R}^{n}\setminus\mathcal{X}])\end{split}$
is the probability that the system (1) starting from $\bm{x}$ will exit the safe set $\mathcal{X}$ at time $t=i$ while stay within $\mathcal{X}$ before $i$, where $i\in\mathbb{N}_{\geq 1}$, we have $\mathbb{P}_{\pi}(S_{\bm{x}})=1-1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{x})-\sum_{i\in\mathbb{N}_{\geq 1}}\mathbb{P}_{\pi}(\wedge_{j=1}^{i-1}[\bm{\phi}_{\pi}^{\bm{x}}(j)\in\mathcal{X}]\wedge[\bm{\phi}_{\pi}^{\bm{x}}(i)\in\mathbb{R}^{n}\setminus\mathcal{X}])=1-V(\bm{x})$. ∎
According to Lemma 1, $V(\bm{x})$ falls within [0,1] for $\bm{x}\in\mathbb{R}^{n}$ and thus it is bounded over $\mathbb{R}^{n}$. We next will show that the value function (2) can be reduced to a bounded solution to a Bellman equation (or, dynamic programming equation) via the dynamic programming principle. A value function characterizes the exact safety probability over finite time horizons and its related dynamic programming equations can be found in [1, 11].
@@
$\begin{split}V(\bm{x})=&1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{x})+1_{\mathcal{X}}(\bm{x})\mathbb{E}_{\pi}[1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{y})+\sum_{i\in\mathbb{N}_{\geq 1}}\prod_{j=0}^{i-1}1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{y}}(j))1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{x}(i))]\\ =&1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{x})+1_{\mathcal{X}}(\bm{x})\mathbb{E}_{\bm{\theta}}[1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{y})+\mathbb{E}_{\pi}[\sum_{i\in\mathbb{N}_{\geq 1}}\prod_{j=0}^{i-1}1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{y}}(j))1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{x}(i))]]\\ =&1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{x})+1_{\mathcal{X}}(\bm{x})\mathbb{E}_{\bm{\theta}}[V(\bm{y})]\\ =&1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{x})+1_{\mathcal{X}}(\bm{x})\mathbb{E}_{\bm{\theta}}[V(\bm{f}(\bm{x},\bm{\theta}))],\\ \end{split}$

### v3

@@ Abstract
This paper investigates necessary and sufficient barrier-like conditions for infinite-horizon safety and reach-avoid verification of stochastic discrete-time systems, derived via a relaxation of the Bellman equations. Unlike prior approaches that primarily focus on sufficient conditions, our work rigorously establishes both necessity and sufficiency for infinite-horizon properties. Safety verification concerns certifying that, starting from a given initial state, the system remains within a safe set at all future time steps with probability at least equal to a specified threshold. For this purpose, we formulate a necessary and sufficient barrier-like condition that captures this infinite-time safety property. In contrast, reach-avoid verification generalizes safety verification by also incorporating reachability. Specifically, it aims to ensure that the probability of the system, starting from a given initial state, eventually reaching a target set while remaining within the safe set until the first hit of the target is no less than a prescribed bound. Under suitable assumptions, we establish two necessary and sufficient barrier-like conditions for this reach-avoid specification.
@@ I Introduction
Temporal verification is crucial in modern systems analysis, particularly in complex systems where temporal behavior is of paramount importance [prajna2007convex]. It involves rigorously examining a system’s adherence to temporal properties, including safety and reach-avoid guarantees, to ensure desired outcomes and avoid undesirable events. Formal methods like model checking [clarke1997model] and theorem proving [manna2012temporal] are indispensable tools in this process, allowing for precise and comprehensive analysis of temporal specifications.
Over the past two decades, barrier certificates have become a powerful tool for safety and reach-avoid verification of dynamical systems. These certificates provide Lyapunov-like guarantees regarding system behavior. The existence of a barrier certificate alone is sufficient to establish the satisfiability of safety and reach-avoid specifications, as demonstrated in [prajna2007convex]. This simplifies the verification process and provides a formal mathematical framework for ensuring the safety and correctness of a system without needing to explicitly evolve it over time. With advances in polynomial optimization, particularly sum-of-squares polynomial optimization, barrier certificates can be computed through convex optimization, especially when the system of interest is polynomial. This further motivates the development of barrier certificate-based methods.
On the other hand, converse theorems for barrier certificates, which focus on the existence of such certificates, have significantly contributed to understanding how safety and reach-avoid criteria can be represented by barrier certificates. These concepts have garnered growing interest since the inception of barrier certificates and have been further investigated in [prajna2005necessity, prajna2007convex, wisniewski2015converse, ratschan2018converse, liu2021converse]. However, there remains a scarcity of research exploring the existence of barrier certificates for stochastic dynamical systems. This work aims to fill this gap.
@@
By relaxing Bellman equations, this paper derives necessary and sufficient barrier-like conditions for verifying safety and reach-avoid properties in stochastic discrete-time systems over infinite-time horizons. The safety verification process involves assessing whether the safety probability that a system, starting from an initial state, will stay within a safe set for all time is greater than or equal to a specified threshold. By relaxing a Bellman equation, one of whose solutions characterizes the exact safety probability, we construct a necessary and sufficient barrier-like condition for safety verification. On the other hand, the reach-avoid verification concerns verifying whether the reach-avoid probability that the system, starting from an initial state, will enter a target set eventually while avoiding unsafe sets before hitting the target, is greater than or equal to a specified threshold. We consider two cases for the reach-avoid verification. In the first case, we assume that, for every state in the safe set but not in the target set, the system will almost surely either reach the target set or exit the safe set in finite time. Under this context, by relaxing a Bellman equation, which possesses a unique bounded solution that characterizes the exact reach-avoid probability, we construct a necessary and sufficient barrier-like condition for the reach-avoid verification. In the second case, we assume that the specified threshold is strictly smaller than the exact reach-avoid probability. Under this context, by relaxing a Bellman equation featuring a unique bounded solution that provides a lower bound of the exact reach-avoid probability, we construct a necessary and sufficient barrier-like condition for the reach-avoid verification.
Related Work
Barrier certificates were initially proposed for deterministic systems as a formal approach to safety verification [prajna2004safety]. Subsequent efforts have focused on adapting and enhancing these functions, broadening their applications [kong2013exponential, ames2019control, ghaffari2018safety, anand2021safety]. However, many real-world applications are subject to stochastic disturbances, prompting the modeling of these systems as stochastic. In the continuous-time stochastic setting, safety verification over the infinite-time horizon using barrier certificates was introduced alongside its deterministic counterpart in [prajna2007framework]. Based on Ville’s Inequality [ville1939etude] and a stopped process, [prajna2007framework] developed a non-negative barrier function and established a sufficient condition for safety verification, certifying upper bounds on the probabilities of entering an unsafe region from specific initial states. This ensures that the system remains within the interior of a state-constrained set until its first encounter with the unsafe set. Building on [kushner1967stochastic], [santoyo2021barrier] formulated a sufficient barrier-like condition for upper-bounding the probability of entering an unsafe region from certain initial states within finite-time frames. The systems in [santoyo2021barrier] include both continuous-time and discrete-time systems. Notably, when the state-constrained set is a robust invariant (i.e., systems initialized within it remain within it under all disturbances) and the unsafe set is a subset of the invariant set, sufficient barrier-like conditions for safety verification of stochastic discrete-time systems were studied in [anand2022k, zhi2024unifying]. Another commonly studied safety property is related to set invariance. This involves justifying lower bounds of safety probabilities, either over an infinite time horizon (i.e., ensuring the system remains within a specified safe set for all time) or finite time horizons (i.e., ensuring the system stays within a given safe set during a specified time period) [abate2008probabilistic]. In other words, it involves justifying upper bounds on the exit probabilities—i.e., the probability that the system will eventually exit a specified safe set or do so within a bounded time horizon. To address this, sufficient barrier-like conditions have been developed for safety verification over both finite-time horizons (e.g., [steinhardt2012finite, laurenti2023unifying, cosner2023robust, cosner2024bounding]) and infinite-time horizons (e.g., [yu2023safe]). Following this, control barrier functions were explored for synthesizing controllers to guarantee safety in [sarkar2020high, wang2021safety]. While finite-time verification suffices for systems with bounded operational horizons, we emphasize the importance of infinite-time safety guarantees—essential for systems requiring robustness against indefinite or unpredictable mission lifetimes. Importantly, the proposed method can also be applied to safety verification over an infinite time horizon, as described in [prajna2007framework].
Regarding reach-avoid verification, a new sufficient barrier-like condition was proposed in [xue2021reach] for the reach-avoid analysis of stochastic discrete-time dynamical systems over an infinite-time horizon. This condition was later extended to stochastic continuous-time dynamical systems in [xue2024]. The condition is constructed by relaxing a set of equations, whose solution characterizes the exact reach-avoid probability of eventually entering a desired target set from an initial state while maintaining safety constraints. In addition, another barrier-like function, called reach-avoid supermartingales, was introduced in [vzikelic2023learning, vzikelic2023compositional] to guarantee reach-avoid specifications and facilitate controller synthesis for stochastic discrete-time systems. This framework assumes that the system evolves within a robust invariant set, with both the unsafe set and target set confined within this invariant domain. However, this strong assumption limits the applicability of the framework, as many systems do not possess compact robust invariant sets, as discussed in [yu2023safe]. These barrier-like conditions aim to lower bound reach-avoid probabilities, as explored in [xue2021reach, vzikelic2023learning, vzikelic2023compositional, xue2024].
@@
This paper is structured as follows: Section II introduces the stochastic discrete-time systems of interest and formulates the safety and reach-avoid verification problems. Section III presents the necessary and sufficient barrier-like conditions for safety verification. Section IV follows with the necessary and sufficient barrier-like conditions for reach-avoid verification. Section V presents two numerical examples that demonstrate the effectiveness of the proposed barrier-like conditions. Finally, Section VI concludes the paper.
@@ II Preliminaries
We start the exposition by a formal introduction of stochastic discrete-time systems and safety/reach-avoid verification problems of interest. Before posing the problem studied, let us introduce some basic notions used throughout this paper: $\mathbb{R}$ denotes the set of real values; $\mathbb{N}$ denotes the set of nonnegative integers; $\mathbb{N}_{\leq k}$ is the set of non-negative integers being less than or equal to $k$; $\mathbb{N}_{\geq k}$ is the set of non-negative integers being larger than or equal to $k$; for sets $\Delta_{1}$ and $\Delta_{2}$, $\Delta_{1}\setminus\Delta_{2}$ denotes the difference of sets $\Delta_{1}$ and $\Delta_{2}$, which is the set of all elements in $\Delta_{1}$ that are not in $\Delta_{2}$; $1_{A}(\bm{x})$ denotes the indicator function in the set $A$, where, if $\bm{x}\in A$, then $1_{A}(\bm{x})=1$ and if $\bm{x}\notin A$, $1_{A}(\bm{x})=0$.
@@ II-A Problem Statement
A disturbance signal $\pi$ is a sample path of the stochastic process $\bm{\theta}(i):\Theta\to\Theta,i\in\mathbb{N}$}, defined on the canonical sample space $\Theta^{\infty}$ equipped with the product topology and Borel $\sigma$-algebra $\mathcal{B}(\Theta^{\infty})$. The probability measure $\mathbb{P}{\pi}:=\mathbb{P}_{\bm{\theta}}^{\infty}$ is the product measure on $\Theta^{\infty}$ induced by the i.i.d. disturbances $\bm{\theta}(0),\bm{\theta}(1),\dots$: $\mathbb{P}_{\pi}=\bigotimes_{i=0}^{\infty}\mathbb{P}_{\bm{\theta}}$, where $\mathbb{P}_{\bm{\theta}}(B)={\rm Prob}(\bm{\theta}(i)\in B)$ for measurable $B\subseteq\Theta$. The expectation $\mathbb{E}_{\pi}[\cdot]$ is defined with respect to $\mathbb{P}_{\pi}$.
@@
The safety and reach-avoid verification for the system (1) over the infinite-time horizon are defined below.
@@
Given a safe set $\mathcal{X}\subseteq\mathbb{R}^{n}$, an initial state $\bm{x}_{0}$, and a lower bound $\epsilon_{1}\in[0,1]$, the safety verification aims to certify that the safety probability $\mathbb{P}_{\pi}(S_{\bm{x}_{0}})$, which denotes the probability that the system (1), starting from the initial state $\bm{x}_{0}$, will stay within the safe set $\mathcal{X}$ for all time, is greater than or equal to $\epsilon_{1}$, i.e.,
$\mathbb{P}_{\pi}(S_{\bm{x}_{0}})\geq\epsilon_{1},$
@@
In the sequel, we will formulate necessary and sufficient barrier-like conditions for certifying $\epsilon_{1}\leq\mathbb{P}_{\pi}(S_{\bm{x}_{0}})$. We note that the method can also be used to construct necessary and sufficient conditions for the safety verification scenario in [prajna2007framework], which involves certifying upper bounds of the probability that the system eventually enters unsafe sets from an initial state while adhering to state-constrained sets. Please refer to Remark 2 in Subsection IV-A. Moreover, under certain assumptions, we will formulate necessary and sufficient barrier-like conditions for certifying $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})\geq\epsilon_{2}$.
@@ III Safety Verification
This section introduces necessary and sufficient barrier-like conditions for certifying lower bounds in safety verification and will detail their construction process. The construction involves constructing and relaxing a Bellman equation, one of whose solutions characterizes the exact safety probability $\mathbb{P}_{\pi}(S_{\bm{x}})$ for $\bm{x}\in\mathbb{R}^{n}$. The Bellman equation is derived from a value function.
We begin by introducing the value function $V(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$, which characterizes the exact safety probability $\mathbb{P}_{\pi}(S_{\bm{x}})$ for each state $\bm{x}\in\mathbb{R}^{n}$,
@@
By definition, $\mathbb{E}_{\pi}[1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(0))]=1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{x})$ holds. Furthermore, since
$\begin{split}\mathbb{E}_{\pi}[\prod_{j=0}^{i-1}1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(i))]=\mathbb{P}_{\pi}(\wedge_{j=0}^{i-1}[\bm{\phi}_{\pi}^{\bm{x}}(j)\in\mathcal{X}]\wedge[\bm{\phi}_{\pi}^{\bm{x}}(i)\in\mathbb{R}^{n}\setminus\mathcal{X}])\end{split}$
is the probability that the system (1) starting from $\bm{x}$ will exit the safe set $\mathcal{X}$ at time $t=i$ while stay within $\mathcal{X}$ before $i$, where $i\in\mathbb{N}_{\geq 1}$, we have
$\begin{split}&\mathbb{E}_{\pi}[1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(0))]+\sum_{i\in\mathbb{N}_{\geq 1}}\mathbb{E}_{\pi}[\prod_{j=0}^{i-1}1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(j))1_{\mathbb{R}^{n}\setminus\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(i))]\\ &=\mathbb{P}_{\pi}(\bm{\phi}_{\pi}^{\bm{x}}(0)\in\mathbb{R}^{n}\setminus\mathcal{X})+\sum_{i\in\mathbb{N}_{\geq 1}}\mathbb{P}_{\pi}(\wedge_{j=0}^{i-1}[\bm{\phi}_{\pi}^{\bm{x}}(j)\in\mathcal{X}]\wedge[\bm{\phi}_{\pi}^{\bm{x}}(i)\in\mathbb{R}^{n}\setminus\mathcal{X}])\\ &=\mathbb{P}_{\pi}(\exists i\in\mathbb{N}.\bm{\phi}_{\pi}^{\bm{x}}(i)\in\mathbb{R}^{n}\setminus\mathcal{X}).\end{split}$
Thus, $\mathbb{E}_{\pi}[g(\bm{x})]=\mathbb{P}_{\pi}(\exists i\in\mathbb{N}.\bm{\phi}_{\pi}^{\bm{x}}(i)\in\mathbb{R}^{n}\setminus\mathcal{X})$. Consequently, $\mathbb{P}_{\pi}(S_{\bm{x}})=1-V(\bm{x})$. ∎
According to Lemma 1, $V(\bm{x})$ falls within [0,1] for $\bm{x}\in\mathbb{R}^{n}$ and thus it is bounded over $\mathbb{R}^{n}$. We next will show that the value function (2) can be reduced to a bounded solution to a Bellman equation (or, dynamic programming equation) via the dynamic programming principle. A value function characterizes the exact safety probability over finite-time horizons and its related dynamic programming equations can be found in [abate2008probabilistic, laurenti2023unifying].
@@
[…]